OpenStudy (anonymous):

A line passes through the point four comma negative one and has a slope of three-halves. If you use this slope to generate another point on the line, what are the coordinates of that point?

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Hi! I remember you from yesterday. Let's recap on what a slope is.

Parth (parthkohli):

The slope is defined as rate of increase of y divided by rate of increase of x

Parth (parthkohli):

So here, the slope is given as 3/4. Meaning that the ratio of increase of y divided by the ratio of increase of x is 3/4. Now, that, in turn means that if x increases by 1, then y increases by 3/4.

Parth (parthkohli):

Do you follow till here?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

So you are given that x = 4 and y = -1 is a solution to this equation.

Parth (parthkohli):

You already know that if you add 1 to x, then you add 3/2 to y.

Parth (parthkohli):

I'm sorry... in the above post I meant 3/2 as well.

Parth (parthkohli):

So, add 1 to x and 3/2 to y. x = 4 y = -1

OpenStudy (anonymous):

1 + 4 = 5, and -1 + 3/2 = 2/4

OpenStudy (anonymous):

And simplifies to 1/2

Parth (parthkohli):

And so, (5,1/2) is a point.

OpenStudy (anonymous):

There isn't a choice that says (5, 1/2

OpenStudy (anonymous):

)

OpenStudy (anonymous):

Nevermind.

Parth (parthkohli):

You have choices?

OpenStudy (anonymous):

I already got the answer (6,2)